OpenStudy (anonymous):

the first 1/5 of a mile is $1.50. For every additional 1/5 of a mile, it cost $0.25. If Marc has $3.75, what is the greatest distance he can travel. p.s. i need an inequality for this problem.

hero (hero):

1.50 + .25x <= 3.75 .25x <= 2.25 x <= 2.25/.25 x<= 9 1/5 + 9/5 = 10/5 = 2 The greatest distance he can travel is 2 miles

hero (hero):

This is a good problem. I can see how it might be difficult to set up.

hero (hero):

Robtobey, I hope you haven't been here this whole time trying to disprove my solution.

OpenStudy (anonymous):

No disapproval.